Operations Bulletin 030417

From GridPP Wiki
Jump to: navigation, search

Bulletin archive

Week commencing Monday 27th March 2017
Task Areas
General updates

  • EGI Operations Broadcast - March 2017
  • HSF Workshop on Visualization software - starting Tuesday March 28, 16hrs - BE Auditorium (6-2-024)

Tuesday 21st March

  • Status of EMI and UMD UI/WN in CVMFS grid.cern.ch ... and planning CentOS7.
  • GridPP VO CVMFS discussion. (Naming?).
  • Minutes of Monday's WLCG ops meeting are available.
  • Reminder of the survey (table) for GridPP38.
  • The final T2 reliability and availability reports for February 2017 are available.
  • Steve asks about ARGO monitoring for the UK and ARC site greyness. File a ticket.
  • Simon points out there is a GridPP DIRAC users list at Imperial connected with support.
  • On 16th: Many ATLAS sites were set in a test mode because of "no SE/SURL protocols defined for output". AGIS had the wrong update values.
  • VAPOR application v2.2 is now online. Map |Faulty Resources | Figures - CPU & Storage split | API.

Tuesday 14th March

  • Ivan: Any experience with adding storage to Fujitsu boxes?
  • IPv6 status and plans. WG | Site status in GridPP.
  • Steve: ARC Brainstorming Camp Summary.
  • The March GDB took place last week at ISGC. Agenda.

Monday 6th March

  • We have received the WLCG T2 R/A figures for February.
    • ALICE
      • All okay
    • ATLAS
      • Oxford 84%:84%
    • CMS
      • All okay
    • LHCb
      • QMUL 84%:84%
      • Durham 88%:88%
    • Site responses:
      • QMUL: No known issues - possibly network related. Following up with LHCb.
      • Durham: Batch system upgrade led to one outage and a University wide internet connection loss led to another.

WLCG Operations Coordination - AgendasWiki Page

Monday 20th March

Monday 6th March

  • There was an ops coordination meeting on Thursday 2nd March - agenda | Minutes.
  • The overview talks were: Mid-SLAs | Machine Job Features.
  • The main summary items:
    • Lessons learned from ATLAS tape usage tests. Follow up via FTS steering group. Will share experience with other experiments.
    • "Mid SLA" resources. During the next months accumulate experience at CERN (providing such resources) and in experiments (using such resources). Review accumulated experience after a couple of months and decide what we need to do in terms of tagging, accounting, etc...
    • MJF deployment is not as successful as desired. Find help (at least temporary) for deployment campaign concentrating on the LHCb sites. See whether this task would become more important considering outcome of the benchmarking discussion.
    • In order to avoid inconsistencies in the naming of the service types along the WLCG IS chain (GocDB, OIM, SAM, experiment-specific systems like Dirac), there will be agreement between representatives from OSG, EGI and GocDB team and IS evolution task force on the policy for introducing of the new service types.
    • The next WLCG ops coordination meeting will be on 6th April.

Monday 27th February

  • The next WLCG Operations Coordination meeting will be Thursday this week starting at 15:30 CET.
  • Themes:
    • SLAs and usage of different kinds of computing resources
    • Machine/Job Features update
  • Experiments, T0, WGs & TFs are requested to add reports to the minutes page at least 1 HOUR before the meeting.
  • Tier-1 and Tier-2 sites are asked to please contact wlcg-ops-coord-chairpeople@cern.ch to have a specific slot added in the agenda if you have non-trivial feedback or announcements to be made.

Tier-1 - Status Page

Tuesday 4th April A reminder that there is a weekly Tier-1 experiment liaison meeting. Notes from the last meeting here

  • Castor SRM updates to version 2.1.16: The LHCb SRMs were successfully updated last Thursday (23rd March). This including upgrading the OS on the systems to SL6. A problem was encountered with the upgrade of the Atlas SRMs on the 28th - with the upgrade being completed the next day. As only two of the four Atlas SRMs were upgraded at first the impact on Atlas was reduced. Remaining SRM upgrades on Thursday (30th).
  • Ongoing work to replace two (of the four) chillers supplying the A/C for the R89 machine room. The first has been replaced and its replacement commissioned.
  • Ongoing tests ongoing with some batch jobs for the LHC VOs running in SL6 containers on worker nodes running SL7.
  • New capacity hardware has been delivered (both CPU and storage). This is now in the racks being cabled up.
  • We have updating our web server. From Wednesday (29th March) www.gridpp.rl.ac.uk points to a new location. References to the old server (lcgwww.gridpp.rl.ac.uk) will no longer work.
  • Out of Hours cover for the CEPH ECHO service is being piloted.
Storage & Data Management - Agendas/Minutes

Wednesday 15 Feb

  • Space tokens! Use them!
  • Leafed through interesting topics (technology permitting) from the CS3 workshop

Wednesday 08 Feb

  • More on accounting (again)

Wednesday 01 Feb

  • Accounting - need to revisit (again); KeyDocs need updating
  • Hardware - Simon's got a disk array to test drive

Wednesday 18 Jan

  • Continued ZFS discussion
  • Updates from the non-LHC VOs, data-wise. Mostly no news, except for DiRAC, Cambridge HPC and Edingburgh EPCC are debugging and testing; LSST expecting to do MC workflows end of Jan.

Tier-2 Evolution - GridPP JIRA

Tuesday 21 March 2017

Tuesday 14 March 2017

Tuesday 28 February 2017

  • ATLAS jobs now running on Birmingham Vac VMs.
  • LHCb is now running its standard DIRAC VMs on Google Compute Engine using GridPP's Vcycle.

Tuesday 21 February 2017

  • Some ALICE production jobs running at Birmingham and Oxford in Vac VMs.

Tuesday 14 February 2017

  • Vac installed at Birmingham. Running LHCb production jobs. Preparing to run ALICE VMs.
  • ALICE VMs switched to using GSI proxies internally rather than host cert/keys.

Accounting - UK Grid Metrics HEPSPEC06 Atlas Dashboard HS06

Monday 16th January

  • The discussion topic for next week will be accounting comparisons. Please note Alessandra's comments last week.

Monday 14th November

  • Alessandra has written an FAQ to extract numbers from ATLAS and APEL avoiding the SSB.

Monday 26th September

  • A problem with the APEL Pub and Sync tests developed last Tuesday and was resolved on Wednesday. This had a temporary impact on the accounting portal.
Documentation - KeyDocs

Tue 7th Mar

I want to get rid of following stale docs...

  • Ten Easy Network Questions
  • GridPP Answers to 10 Easy Network Questions
  • SL5 status
  • Glasgow Middleware Operations Logbook
  • London Tier2 Shares
  • 23rd June 2010 Special Topic: NGS Technical Roadmap - Pete Gronbech
  • SAM availability: Monthly summary table
  • Transfer Tests Birmingham Endpoints
  • EVO Tips and Tricks
  • BaBar: bbrbsub270

Monday 13th Feb

  • Please check the keydocs page. Would all owners please review their pages and prepare a short note on how they will bring things up-to-date in the coming months. This will be discussed at a core ops meeting (poll to be circulated shortly).
  • There are several EGI operations document updates out for review. Please make comments to Jeremy by 20th Febraury for inclusion in the UK feedback:

Monday 30th January

The fusion VO has been removed from GridPP Approved VOs.

Monday 9th January

  • Tom put these together:
    • A snapshot of the GridPP UserGuide is now available as an offline document in the Zenodo repository:
    • ...and the v1.0 release of the LaTeX code used to compile it is here.
    • There are some minor differences (mainly hyperlink-related) to the online version [1], which has also had a v1.0 release here.

Tue 22nd Nov

  • More on Accounting audit tools

The process for ARC, Toque and VAC is now documented here.


General note

See the worst KeyDocs list for documents needing review now and the names of the responsible people.

Interoperation - EGI ops agendas

Tuesday 21st March

Monday 13th February

  • There was an ops meeting on 13th February. Agenda
    • frontier-squid-3 in UMD 4.4.0 - changes to config
    • Assess IPv6 readiness:
      • Resource Centres: assess the IPv6 readiness of the site infrastructure (real machines, cloud managers)
      • NGIs/ROCs please start discussing with sites and provide suggestions for the overall plan
    • Decommissioning of dCache 2.10
    • Discussed stale ARGO results; being discussed with Product Teams.

Monday 16th January

  • There was an ops meeting on 9th January. Agenda.

Monday 7th November

  • There was an EGI Ops meeting today: agenda
  • UMD 3.14.5 released today
    • VOMS 3.5.0, which makes RFC proxies the default for voms-proxy-init
  • UMD 4.3.0 'October' release, release candidate ready, to be released by end of this week, including:
    • ARC, GFAL2, XROOT, Davix, dCache, ARGUS, Gridsite, edg-mkgrid, umd-release for CentOS7
  • please start using UMD4/SL6 or UMD4/CentOS7 instead of UMD3/SL6 & please don't use anymore EMI3
    • (think there may be a campaign around this soon)
  • Downtimes due to the vulnerability CVE-2016-5195: request an A/R recomputation
    • All the resource centres that were affected by the vulnerability CVE-2016-5195 and that declared a downtime between 2016-10-20 16:00 UTC and 2016-10-31 18:00 UTC are invited to request a recomputation of A/R figures for the days in which the downtime was ongoing.
Monitoring - Links MyWLCG

Monday 13th Febraury

  • This category is pretty much inactive. Are there any topics under "monitoring" that anyone wants reported at this ops meeting? If not we will remove this section from the regular updates area of the bulletin and just leave the main links.

Tuesday 1st December

Tuesday 16th June

  • F Melaccio & D Crooks decided to add a FAQs section devoted to common monitoring issues under the monitoring page.
  • Feedback welcome.

On-duty - Dashboard ROD rota

20th March

  • Another quiet week. One ticket to Oxford and a new availablility ticket to Glasgow. Other than that, most of the tickets have been there for a couple of weeks.

27th February

  • Forwarded two tickets (RALPP and ECDF) to the monitoring team, as this looks like more ARC shenanigans.
  • There's also a bunch of low availability tickets which clog up the system.

Tuesday 21st February

  • Very quiet!
  • Rota responses enable planning through to May.

Monday 16th January

  • Rota to be updated this week. Gordon is on shift this week.
Rollout Status WLCG Baseline

Monday 13th February

  • Please can we hear something from Raul and Sam (or others contributing to the middleware readiness work) about news in this area at the ops meeting next week? Thank you.
  • Does anyone want specific questions in this area tackled in a roundtable discussion?

Tuesday 7th December

  • Raul reports: validation of site BDII on Centos 7 done.

Tuesday 15th September

Tuesday 12th May

  • MW Readiness WG meeting Wed May 6th at 4pm. Attended by Raul, Matt, Sam and Jeremy.


Security - Incident Procedure Policies Rota

Tuesday 28th March

  • The following advisories were issued under [TLP:AMBER], check the advisory for details and if you might be affected -
    • 1 High risk vulnerability advisory issued by EGI SVG potentially affecting a number of services.
    • 1 Critical risk vulnerability advisory issued by EGI SVG affecting 1 service.
    • 1 Security Incident reported via the EGI CSIRT, not in the UK.
  • Updated IGTF distribution 1.82 available - updated CA root certificates [1]
  • WISE - Security for collaborating infrastructures meeting this week [2]

Tuesday 21st March

  • One advisory to follow-up.

Tuesday 14th March

  • EGI SVG Advisory [TLP:WHITE] 'HIGH' risk linux kernel privilege escalation vulnerability CVE-2017-2636 [EGI-SVG-CVE-2017-2636]
  • One incident reported by one UK site.
  • UK Security Team meeting this afternoon.

Services - PerfSonar production dashboard |PerfSonar development dashboard | GridPP VOMS

- This includes notifying of (inter)national services that will have an outage in the coming weeks or will be impacted by work elsewhere. (Cross-check the Tier-1 update).

Monday 6th March

  • Last week it was noted that some ATLAS RIPE probes (the USB ones) have had hardware errors. If you need a replacement and RIPE are slow to respond then contact Jeremy.
  • Is anyone running any tests with their RIPE probe?

Tuesday 14th February

  • The development dashboard is now operational again. However the production dashboard is not currently displaying latency results (OSG ticket submitted).
  • Birmingham pS host had lost its entire network connection and both Lancaster's hosts their IPv6 connection. Durham results are incomplete, GGUS ticket submitted (#126587).

Tuesday 25th October

  • Duncan has recreated the UK perfSONAR mesh. Link here!


Monday 27th March 2017, 15.15 BST
26 Open UK Tickets this week.

Nearly a third of the UK tickets are not On Hold but not received an update in over 10 days. Before I go over all the tickets next week please could everyone check their older tickets. In light of that I'll have a delicate review of the tickets this week.

126184 (26/1)
Has everyone who had input on the atlas site monitoring survey said their piece? If yes then this ticket has done its job. In progress (22/3)

126956 (6/3)
I cheekily set this Liverpool availability ticket on hold as per the S.O.P. On Hold (27/3)

I did similar to this just as unjust Brunel ticket: 127126

126976 (6/3)
This Sno+ ticket looks to be solved (it was fixed by adjusting the acls and directory permissions after implementing the new spacetoken) - so the ticket can probably be closed. In Progress (25/3)

Tools - MyEGI Nagios

29th November 2016

LSST monitoring is available through vo-nagios


13th September 2016

VOs - GridPP VOMS VO IDs Approved VO table
Site Updates


Meeting Summaries
Project Management Board - MembersMinutes Quarterly Reports


GridPP ops meeting - Agendas Actions Core Tasks


RAL Tier-1 Experiment Liaison Meeting (Wednesday 13:30) Agenda Meeting takes place on Vidyo.

Highlights from this meeting are now included in the Tier1 report farther up this page.

WLCG Grid Deployment Board - Agendas MB agendas


NGI UK - Homepage CA


UK ATLAS - Shifter view News & Links

Atlas S&C week 2-6 Feb 2015


• Prodsys-2 in production since Dec 1st

• Deployment has not been transparent , many issued has been solved, the grid is filled again

• MC15 is expected to start soon, waiting for physics validations, evgen testing is underway and close to finalised.. Simulation expected to be broadly similar to MC14, no blockers expected.


• Rucio in production since Dec 1st and is ready for LHC RUN-2. Some fields need improvements, including transfer and deletion agents, documentation and monitoring.

Rucio dumps available.

Dark data cleaning

files declaration . Only Only DDM ops can issue lost files declaration for now, cloud support needs to fill a ticket.

• Webdav panda functional tests with Hammercloud are ongoing


Main page

DDM Accounting




• ASAP (ATLAS Site Availability Performance) in place. Every 3 months the T2s sites performing BELOW 80% are reported to the International Computing Board.





  • N/A
To note

  • N/A